OpenStudy (anonymous):

the product of the digits of a two digit number is 24.If the units's digit exceeds twice its ten's digit by 2: find the number

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

help?

ganeshie8 (ganeshie8):

say the number = ab

ganeshie8 (ganeshie8):

\[a\times b = 24\tag{1}\] \[b = 2a+2\tag{2}\] solve

OpenStudy (anonymous):

???

OpenStudy (anonymous):

its a quadratic equation problem

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

did u get how ganeshie8 got those 2 equation?

OpenStudy (anonymous):

its quadratic....

ganeshie8 (ganeshie8):

they are system of equations

ganeshie8 (ganeshie8):

you can get a simple quadratic by combining them

hartnn (hartnn):

if you plug in 'b' in the first equation,you will get a quadratic in a

OpenStudy (anonymous):

im doing this topic after 1 yr soo im confused

hartnn (hartnn):

2 digit number : ab the product of the digits of a two digit number = \(a\times b\) is 24 : \(a\times b =24\) good so far ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

if the units's digit exceeds twice its ten's digit by 2 b= 2a+2 this?

hartnn (hartnn):

units digit is b tens digit is a

OpenStudy (anonymous):

yes so now what??

hartnn (hartnn):

plug in b =2a+2 in ab =24

OpenStudy (anonymous):

????

hartnn (hartnn):